# {cutout}00.3cm22cm1 empty    PointMamba: A Simple State Space Model for Point Cloud Analysis

### v2

@@ Abstract
00footnotetext: Dingkang Liang, Xin Zhou, and Xinyu Wang make equal contributions.
@@ 1 Introduction
Point cloud analysis is one of the fundamental tasks in computer vision and has a wide range of real-world applications [48; 31; 65; 5]. It is a challenging task due to the intrinsic irregularity and sparsity of point clouds. To deal with the issues, there has been rapid progress in deep learning-based methods [37; 29; 40; 42], consistently pushing the performance to the new record.
Recently, transformers have shown great potential in point cloud analysis. The key to the transformer is the attention mechanism, which can effectively capture the relationship of a set of points. By further combining self-supervised learning paradigms and fine-tuning the downstream tasks, these transformer-based methods [63; 29; 41] have obtained superior performance. For example, Point-BERT [63] generates discrete point tokens containing meaningful local information from masked tokens. PointMAE [29] follows the pipeline of ViT-based MAE [21], reconstructing the masked point patches. ReCon [41] learn through ensemble distillation from both generative modeling teachers and single/cross-modal contrastive teachers.
@@
We note the recent advance of the state space models (SSMs) [12; 54; 24]. Specifically, the Structured State Space Sequence Model [15] (S4) has emerged as a promising class of architectures for sequence modeling. Based on S4, Mamba [12] adopt time-varying parameters to the SSM, proposing an efficient hardware-aware algorithm, having global receptive fields with linear complexity. Currently, those SSM-based models mainly focus on long-range and casual data, e.g., language understanding [35], content-based reasoning [12]. Recently, a few concurrent methods [69; 30] pay attention to 2D vision tasks (e.g., classification, segmentation).
Inspired by the successes of SSM in NLP, this paper aims to explore the potential of SSM in point cloud analysis tasks. Through pilot experiments, we find that simply using Mamba [12] can not achieve ideal performance. We argue that the challenge might be the unidirectional modeling of default SSM. In contrast, the self-attention of the transformer is invariant to the permutation and cardinality of the input elements. Thus, in this paper, we introduce the Point State Space Model (denoted as PointMamba), which has global modeling with linear complexity. Specifically, following the standard transformer-based methods [29; 63], we first adopt a simple point tokenizer to generate a series of point tokens. Then, a simple reordering strategy is proposed to scan data along a specific order, enabling the model to causally capture the point cloud structure. Finally, the reordered point tokens will be fed into the mamba encoder to implement global modeling.
@@ 2.1 Point Cloud Transformers
Transformer [51] can capture global input dependencies using the Attention mechanism and has been dominant in NLP [8; 44; 45; 1] and 2D Vision [10; 32; 2; 28]. In 3D vision, the Transformer architecture continues to evolve. Point-BERT [63] and Point-MAE [37] introduce a standard Transformer architecture for self-supervised learning and is applicable to different downstream tasks, as demonstrated in several works [9; 41; 4; 64]. Other researchers [18; 68; 55] mainly concentrate on enhancing Transformers for point clouds. The PCT [18] conducts global attention directly on the point cloud. Point Transformer [68] applies vector attention theory [67] to perform local attention between each point and its adjacent points. The later Point Transformer series [59; 58] further extends the performance and efficiency of the Transformer for different tasks.
Standard transformers can be smoothly integrated into autoencoders using an encoder-decoder design [21], which makes this structure ideal for pre-training and leads to significant performance improvements in downstream tasks [63; 37]. However, the Attention mechanism, a crucial module of the Transformer, has a time complexity of $O(n^{2}d)$, where $n$ represents the length of the input token sequence and $d$ represents the dimension of the Transformer. This implies that as the input sequence grows, the operational efficiency of the Transformer is significantly constrained.
@@ 2.2 State Space Models
Drawing inspiration from control theory, linear state space equations have been combined with deep learning to address sequential data modeling. Initiated by a HiPPO matrix [13], LSSL [16] utilizes linear state space equations for sequence data modeling. Subsequently, the Structured State Space Sequence Model [15] (S4) employs a linear state space for contextualization and shows strong performance on various sequence modeling tasks,especially with lengthy sequences. To avoid high computational burden, HTTYH [17], DSS [19] and S4D [14] find that a diagonal matrix in S4 can replace the Diagonal Plus Low Rank structure without compromising performance. SGConv [27] also provides an alternative approach for treating S4 as a global conventional model. To make S4 faster, GSS [35] leverages the gating structure in the gated attention unit to reduce the dimension of the state space module. Most importantly, the S5 layer [49] introduces the parallel scan and the MIMO SSM, enabling the state space model to be efficiently utilized and widely implemented. As ViS4mer [24], S4ND [36], TranS4mer [25] and The Selective S4 model [54] have applied the state space model in the field of computer vision, achieving preliminary success.
Recently, Mamba [12] has achieved a groundbreaking milestone with its linear-time inference and efficient training process by integrating selection mechanism and hardware-aware algorithms into previous works [14; 19; 35]. Building on the success of Mamba, MoE-Mamba [38] combines Mixture of Experts with Mamba to unlock the potential of SSMs for scaling, achieving Transformer-like performance. Operating on bytes, MambaByte [53] outperforms other byte-level models over several datasets and shows competitive results with subword Transformers. Graph-Mamba [52] enhances long-range context modeling in graph networks by integrating a Mamba block with the input-dependent node selection mechanism. For visual application, Vision Mamba [30] and VMamba [69] use bidirectional Vim Block and the Cross-Scan Module respectively to gain data-dependent global visual context. At the same time, U-Mamba [33] and other works [60; 47] show superior performance in medical image segmentation. These are concurrent works that have appeared while this manuscript is being prepared. However, the application of the Mamba for point cloud analysis remains unexplored. In this work, we delve into the potential of Mamba in 3D vision and propose PointMamba.
@@ 3.2.2 Point Tokenizer
Building on the pioneer works [63; 37], we utilize the Farthest Point Sampling (FPS) and K-Nearest Neighborhood (KNN) algorithm to partition the input point cloud into irregular point patches. Specifically, given an input point cloud with $M$ points $I\in\mathbb{R}^{M\times 3}$, the FPS is applied to sample $n$ key points from the original point cloud $I$. Subsequently, the KNN algorithm selects $k$ nearest neighbors for each key point, forming $n$ patches $P\in\mathbb{R}^{n\times k\times 3}$ with patch size $k$. To aggregate local information, points within each patch are normalized by subtracting the key point to obtain relative coordinates. We finally map the unbiased local patches to feature space using a lightweight PointNet [39], obtaining point tokens $E_{0}\in\mathbb{R}^{n\times C}$.
@@ 3.2.3 Reordering Strategy
A simple approach is to scan data along a specific order, enabling the model to causally capture the structure of point cloud. As shown in Fig. 2, the original random sequence used in transformer-based counterparts [63; 37] lacks a coherent geometric order, making it challenging for unidirectional modeling Mamba. To address this, we arrange point tokens $E_{0}\in\mathbb{R}^{n\times C}$ based on the geometric coordinates of its clustering center, sorting and concatenating along the $x$, $y$, and $z$ axes, thereby tripling the length of the point token (i.e., $E_{0}^{\prime}\in\mathbb{R}^{3n\times C}$). We argue that applying the reordering strategy improves Mamba’s global modeling capability by providing a more logical geometric scanning order. We find that despite increasing the sequence length, the reordering strategy results in a considerable performance improvement without a significant increase in FLOPs.
@@ 4 Experiments
In this section, we first present the implementation details. Then, we evaluate the proposed model with two downstream tasks, including real-world object classification and part segmentation. We also carry on an ablation study for our PointMamba.
@@ 4.1 Implementation Details
(ShapeNetCore) (ScanObjectNN) (ShapeNetPart)
@@
Learning rate 1e-3 5e-4 2e-4
Weight decay 5e-2 5e-2 5e-2
Learning rate scheduler cosine cosine cosine
Training epochs 300 300 300
Warmup epochs 10 10 10
Batch size 128 32 16
Num. of encoder layers $N$ 12 12 12
Num. of decoder layers 4 - -
Input points $M$ 1024 2048 2048
Num. of patches $n$ 64 128 128
Patch size $k$ 32 32 32
Augmentation Scale&Trans Scale&Trans -
SSM is a newcomer to 3D point cloud analysis, and no existing works currently detail the specific implementation. For pre-training, we follow Point-MAE [37], using the ShapeNetCore training set, a subset of ShapeNet [3]. When presenting the results for downstream tasks, we include both scratch and pre-training scenarios. All experiments are carried out using a single RTX 4090 GPU. The implementation details can be found in Tab. 1.
Table 2: Object classification on ScanObjectNN [50]. We evaluate PointMamba on three variants, with PB-T50-RS being the most challenging. Accuracy (%) is reported. $*{}^{*}$ denotes reproduced results.
@@
Training from pre-training, following Point-MAE [37]
@@
Table 3: Part segmentation on the ShapeNetPart [62]. The mIoU for all classes (Cls.) and for all instances (Inst.) are reported. $*{}^{*}$ denotes reproduced results.
@@
Training from pre-training, following Point-MAE [37]
@@ 4.2 Compared with Transformer-based Counterpart
Real-World Object Classification on ScanObjectNN. ScanObjectNN [50] is a highly challenging 3D dataset consisting of about 15,000 objects across 15 categories. These objects are scanned from real-world indoor scenes with cluttered backgrounds, adding to the complexity of object classification. Unlike ViT [10], which employs an attention mechanism for global interactions, we do not employ the classification token $[\mathtt{CLS}]$ as PointMamba can only perform unidirectional modeling. Instead, we simply utilize the average pooling of the last block’s output for classification. In Tab. 2, we conduct experiments on three versions of ScanObjectNN (i.e., OBJ-BG, OBJ-ONLY, and PB-T50-RS) trained from scratch and pre-training, each with increasing complexity. It is worth noting that our PointMamba outperforms the Point-MAE in all three variants when training from scratch while reducing 44.3% parameters and 25% FLOPs. Specifically, compared with reproduced results, PointMamba achieves increases of 1.55%, 0.86%, and 1.7% in the three ScanObjectNN variants. Even when trained from pre-training, our model still achieves competitive results. Specifically, the PointMamba outperforms Point-BERT 3.28%, 0.35% and 1.18% on three variants, exceeding Point-MAE 0.69% and 0.18% on OBJ-BG and OBJ-ONLY. Another thing to notice is that, with the same pre-training strategy, the improvement of our PointMamba is not as significant as Point-MAE but still surpasses Point-MAE in most cases, suggesting that a customized pre-training strategy for PointMamba is a worthwhile research direction, which left our future work.
@@
Part Segmentation on ShapeNetPart. Part segmentation is another challenging task aiming to predict a more detailed label for each point within a sample. We assess the effectiveness of PointMamba on ShapeNetPart [62], which includes 16 typical real-life shape categories and 16,881 samples. We collect 2048 points in 128 point patches, then adopt the reordering strategy to triple the input. To ensure fair comparisons, our segmentation head follows Point-MAE [37] and utilizes learned features from the 3-rd, 7-th, and 11-th layers of the Mamba block. We combine the three levels of features and then use max and average pooling to obtain global features. After concatenating per-point features with the global features, we use a Linear to predict the label for each point. As shown in Tab. 3, we report mean IoU (mIoU) for all classes (Cls.) and all instances (Inst.). Our PointMamba achieve competitive results while significantly reducing parameters. Note that the increase in parameters and FLOPs mainly comes from the head part. Moreover, we present qualitative results of our PointMamba and Point-MAE in Fig. 3. It can be found that PointMamba achieves highly competitive results to the Transformer-base counterpart [29].
@@ 4.3 Ablation study
Table 4: The effect of reordering strategy. $*{}^{*}$ denotes reproduced results.
@@
Point-MAE ($*{}^{*}$ 22.1 4.8 86.75 86.92
Point-MAE ($*{}^{*}$ 22.1 10.2 86.40 86.57
Point-MAE ($*{}^{*}$ 22.1 18.4 86.06 86.23
@@
Analysis on classification token. Previous works [10; 63; 37] often use a classification token $[\mathtt{CLS}]$ as a global token for classification. As in Tab. 5, we find that without $[\mathtt{CLS}]$ and utilizing only the average pooling of the final block’s output yields the best results for PointMamba.
@@
Analysis on dimension $C$. The dimension $C$ for Mamba block is a crucial hyper-parameter for controlling the model’s scale. As in Tab. 6, we empirically sets $C$ to 384 by considering the trade-off in performance across three variants.

### v3

@@ Abstract
00footnotetext: Dingkang Liang, Xin Zhou, and Xinyu Wang make equal contribution.
@@ 1 Introduction
Point cloud analysis is one of the fundamental tasks in computer vision and has a wide range of real-world applications [48; 31; 66; 5]. It is a challenging task due to the intrinsic irregularity and sparsity of point clouds. To deal with the issues, there has been rapid progress in deep learning-based methods [37; 29; 41; 43], consistently pushing the performance to the new record.
Recently, transformers have shown great potential in point cloud analysis. The key to the transformer is the attention mechanism, which can effectively capture the relationship of a set of points. By further combining self-supervised learning paradigms and fine-tuning the downstream tasks, these transformer-based methods [64; 29; 42] have obtained superior performance. For example, Point-BERT [64] generates discrete point tokens containing meaningful local information from masked tokens. PointMAE [29] follows the pipeline of ViT-based MAE [21], reconstructing the masked point patches. ReCon [42] learn through ensemble distillation from both generative modeling teachers and single/cross-modal contrastive teachers.
@@
We note the recent advance of the state space models (SSMs) [12; 54; 24]. Specifically, the Structured State Space Sequence Model [15] (S4) has emerged as a promising class of architectures for sequence modeling. Based on S4, Mamba [12] adopt time-varying parameters to the SSM, proposing an efficient hardware-aware algorithm, having global receptive fields with linear complexity. Currently, those SSM-based models mainly focus on long-range and casual data, e.g., language understanding [35], content-based reasoning [12]. Recently, a few concurrent methods [70; 30] pay attention to 2D vision tasks (e.g., classification, segmentation).
Inspired by the successes of SSM in NLP, this paper aims to explore the potential of SSM in point cloud analysis tasks. Through pilot experiments, we find that simply using Mamba [12] can not achieve ideal performance. We argue that the challenge might be the unidirectional modeling of default SSM. In contrast, the self-attention of the transformer is invariant to the permutation and cardinality of the input elements. Thus, in this paper, we introduce the Point State Space Model (denoted as PointMamba), which has global modeling with linear complexity. Specifically, following the standard transformer-based methods [29; 64], we first adopt a simple point tokenizer to generate a series of point tokens. Then, a simple reordering strategy is proposed to scan data along a specific order, enabling the model to causally capture the point cloud structure. Finally, the reordered point tokens will be fed into the mamba encoder to implement global modeling.
@@ 2.1 Point Cloud Transformers
Transformer [51] can capture global input dependencies using the Attention mechanism and has been dominant in NLP [8; 44; 45; 1] and 2D Vision [10; 32; 2; 28]. In 3D vision, the Transformer architecture continues to evolve. Point-BERT [64] and Point-MAE [37] introduce a standard Transformer architecture for self-supervised learning and is applicable to different downstream tasks, as demonstrated in several works [9; 42; 4; 65]. Other researchers [18; 69; 55] mainly concentrate on enhancing Transformers for point clouds. The PCT [18] conducts global attention directly on the point cloud. Point Transformer [69] applies vector attention theory [68] to perform local attention between each point and its adjacent points. The later Point Transformer series [59; 58] further extends the performance and efficiency of the Transformer for different tasks.
Standard transformers can be smoothly integrated into autoencoders using an encoder-decoder design [21], which makes this structure ideal for pre-training and leads to significant performance improvements in downstream tasks [64; 37]. However, the Attention mechanism, a crucial module of the Transformer, has a time complexity of $O(n^{2}d)$, where $n$ represents the length of the input token sequence and $d$ represents the dimension of the Transformer. This implies that as the input sequence grows, the operational efficiency of the Transformer is significantly constrained.
@@ 2.2 State Space Models
Drawing inspiration from control theory, linear state space equations have been combined with deep learning to address sequential data modeling. Initiated by a HiPPO matrix [13], LSSL [16] utilizes linear state space equations for sequence data modeling. Subsequently, the Structured State Space Sequence Model [15] (S4) employs a linear state space for contextualization and shows strong performance on various sequence modeling tasks, especially with lengthy sequences. To avoid high computational burden, HTTYH [17], DSS [19], and S4D [14] find that a diagonal matrix in S4 can replace the Diagonal Plus Low Rank structure without compromising performance. SGConv [27] also provides an alternative approach for treating S4 as a global conventional model. To make S4 faster, GSS [35] leverages the gating structure in the gated attention unit to reduce the dimension of the state space module. Most importantly, the S5 layer [49] introduces the parallel scan and the MIMO SSM, enabling the state space model to be efficiently utilized and widely implemented. As ViS4mer [24], S4ND [36], TranS4mer [25] and The Selective S4 model [54] have applied the state space model in the field of computer vision, achieving preliminary success.
Recently, Mamba [12] has achieved a groundbreaking milestone with its linear-time inference and efficient training process by integrating selection mechanism and hardware-aware algorithms into previous works [14; 19; 35]. Building on the success of Mamba, MoE-Mamba [39] combines Mixture of Experts with Mamba to unlock the potential of SSMs for scaling, achieving Transformer-like performance. Operating on bytes, MambaByte [53] outperforms other byte-level models over several datasets and shows competitive results with subword Transformers. Graph-Mamba [52] enhances long-range context modeling in graph networks by integrating a Mamba block with the input-dependent node selection mechanism. For visual application, Vision Mamba [30] and VMamba [70] use bidirectional Vim Block and the Cross-Scan Module respectively to gain data-dependent global visual context. At the same time, U-Mamba [33] and other works [61; 47] show superior performance in medical image segmentation. These are concurrent works that have appeared while this manuscript is being prepared. However, the application of the Mamba for point cloud analysis remains unexplored. In this work, we delve into the potential of Mamba in 3D vision and propose PointMamba.
@@ 3.2.2 Point Tokenizer
Building on the pioneer works [64; 37], we utilize the Farthest Point Sampling (FPS) and K-Nearest Neighborhood (KNN) algorithm to partition the input point cloud into irregular point patches. Specifically, given an input point cloud with $M$ points $I\in\mathbb{R}^{M\times 3}$, the FPS is applied to sample $n$ key points from the original point cloud $I$. Subsequently, the KNN algorithm selects $k$ nearest neighbors for each key point, forming $n$ patches $P\in\mathbb{R}^{n\times k\times 3}$ with patch size $k$. To aggregate local information, points within each patch are normalized by subtracting the key point to obtain relative coordinates. We finally map the unbiased local patches to feature space using a lightweight PointNet [40], obtaining point tokens $E_{0}\in\mathbb{R}^{n\times C}$.
@@ 3.2.3 Reordering Strategy
A simple approach is to scan data along a specific order, enabling the model to causally capture the structure of the point cloud. As shown in Fig. 2, the original random sequence used in transformer-based counterparts [64; 37] lacks a coherent geometric order, making it challenging for unidirectional modeling Mamba. To address this, we arrange point tokens $E_{0}\in\mathbb{R}^{n\times C}$ based on the geometric coordinates of its clustering center, sorting and concatenating along the $x$, $y$, and $z$ axes, thereby tripling the length of the point token (i.e., $E_{0}^{\prime}\in\mathbb{R}^{3n\times C}$). We argue that applying the reordering strategy improves Mamba’s global modeling capability by providing a more logical geometric scanning order. We find that despite increasing the sequence length, the reordering strategy results in a considerable performance improvement without a significant increase in FLOPs.
@@ 4 Experiments
In this section, we first present the implementation details. Then, we evaluate the proposed model with three downstream tasks, including synthetic object classification, real-world object classification and part segmentation. We also carry on an ablation study for our PointMamba.
@@ 4.1 Implementation Details
ShapeNetCore ModelNet40 ScanObjectNN ShapeNetPart
@@
Learning rate 1e-3 3e-4 5e-4 2e-4
Weight decay 5e-2 5e-2 5e-2 5e-2
Learning rate scheduler cosine cosine cosine cosine
Training epochs 300 300 300 300
Warmup epochs 10 10 10 10
Batch size 128 32 32 16
Num. of encoder layers $N$ 12 12 12 12
Num. of decoder layers 4 - - -
Input points $M$ 1024 1024 2048 2048
Num. of patches $n$ 64 64 128 128
Patch size $k$ 32 32 32 32
Augmentation Scale&Trans Scale&Trans Scale&Trans -
SSM is a newcomer to 3D point cloud analysis, and no existing works currently detail the specific implementation. For pre-training, we follow Point-MAE [37], using the ShapeNetCore training set, a subset of ShapeNet [3]. When presenting the results for downstream tasks, we include both scratch and pre-training scenarios.
For classification tasks, we do not use the classification token [CLS] as PointMamba can only perform unidirectional modeling, unlike ViT[10], which utilizes an attention mechanism for global interactions. Instead, we simply select the average pooling of the last block’s output for classification. For part segmentation, our segmentation head follows Point-MAE [37] and leverages features from the 3-rd, 7-th, and 11-th layers of the Mamba block for fair comparisons. We merge the three feature levels, apply max and average pooling for global features, and then concatenate per-point features with the global ones. Finally, a Linear layer is used to forecast the label for each point.
All experiments are carried out using a single RTX 4090 GPU. More implementation details can be found in Tab. 1.
Table 2: Object classification on ModelNet40 [60]. Overall accuracy (%) is reported. denotes reproduced results. The results obtained from inputting 1024 points without voting are reported.
Methods Backbone Param. (M) $\downarrow$ FLOPs (G) $\downarrow$ Points Num. OA (%) $\uparrow$
PointNet [40] - 3.5 0.5 1k 89.2
PointNet++ [41] - 1.5 1.7 1k 90.7
PointCNN [26] - 0.6 - 1k 92.2
DGCNN [38] - 1.8 2.4 1k 92.9
PointNeXt [43] - 1.4 1.6 1k 92.9
PCT [18] Transformer-based 2.9 2.3 1k 93.2
OctFormer [55] Transformer-based - - 1k 92.7
Training from scratch
Point-MAE [37] Transformer-based 22.1 2.4 1k 92.3
PointMamba (ours) Mamba-based 12.3 1.8 1k 92.4
Training from pre-training
Point-BERT [64] Transformer-based 22.1 2.4 1k 92.7
Point-MAE [37] Transformer-based 22.1 2.4 1k 93.2
PointMamba (ours) Mamba-based 12.3 1.8 1k 93.6
Table 3: Object classification on ScanObjectNN [50]. We evaluate PointMamba on three variants, with PB-T50-RS being the most challenging. Accuracy (%) is reported. denotes reproduced results, indicates that using simple rotational augmentation [9] for training.
@@
Training from pre-training
@@
PointMamba (ours) Mamba-based 12.3 3.6 93.29 91.91 88.17
Table 4: Part segmentation on the ShapeNetPart [63]. The mIoU for all classes (Cls.) and for all instances (Inst.) are reported. denotes reproduced results.
@@
Training from pre-training
@@ 4.2 Compared with Transformer-based Counterpart
Synthetic Object Classification on ModelNet40. ModelNet40 [60] is a pristine 3D CAD dataset comprising 40 categories. As shown in Tab. 2, we report overall Accuracy without voting. PointMamba achieves competitive results compared to Point-MAE from scratch while using fewer parameters and FLOPs. Using the MAE-like pre-training strategy, our PointMamba achieved a 1.2% performance improvement, outperforming Point-BERT and Point-MAE by 0.9% and 0.4%, respectively. The pre-trained PointMamba can also achieve competitive results against some point cloud-specific transformer models such as PCT and OctFormer.
Real-World Object Classification on ScanObjectNN. ScanObjectNN [50] is a highly challenging 3D dataset consisting of about 15,000 objects across 15 categories. These objects are scanned from real-world indoor scenes with cluttered backgrounds, adding to the complexity of object classification. In Tab. 3, we conduct experiments on three versions of ScanObjectNN (i.e., OBJ-BG, OBJ-ONLY, and PB-T50-RS) trained from scratch and pre-training, each with increasing complexity. It is worth noting that our PointMamba outperforms the Point-MAE in all three variants when training from scratch while reducing 44.3% parameters and 25% FLOPs. Specifically, compared with reproduced results, PointMamba achieves increases of 1.55%, 0.86%, and 1.7% in the three ScanObjectNN variants. Even when trained from pre-training, our model still achieves competitive results. Specifically, the PointMamba outperforms Point-BERT 3.28%, 0.35% and 1.18% on three variants, exceeding Point-MAE 0.69% and 0.18% on OBJ-BG and OBJ-ONLY. Further adopting simple rotational augmentation [9] can further improve the model performance, showcasing the great potential for real-world classification. Another thing to notice is that, with the same pre-training strategy, the improvement of our PointMamba is not as significant as Point-MAE but still surpasses Point-MAE in most cases, suggesting that a customized pre-training strategy for PointMamba is a worthwhile research direction, which left our future work.
@@
Part Segmentation on ShapeNetPart. Part segmentation on ShapeNetPart [63] is another challenging task aiming to predict a more detailed label for each point within a sample. We collect 2048 points in 128 point patches, then adopt the reordering strategy to triple the input. As shown in Tab. 4, we report mean IoU (mIoU) for all classes (Cls.) and all instances (Inst.). Our PointMamba achieve competitive results while significantly reducing parameters. Note that the increase in parameters and FLOPs in segmentation mainly comes from the head part. We do not use the U-Net structure for fair comparison, which may limit the performance. Moreover, we present qualitative results of our PointMamba and Point-MAE in Fig. 3. It can be found that PointMamba achieves highly competitive results to the Transformer-base counterpart [29].
@@ 4.3 Ablation study
Table 5: The effect of reordering strategy. denotes reproduced results.
@@
Point-MAE (1x) 22.1 4.8 86.75 86.92
Point-MAE (3x) 22.1 10.2 86.40 86.57
Point-MAE (6x) 22.1 18.4 86.06 86.23
@@
Analysis on classification token. Previous works [10; 64; 37] often use a classification token $[\mathtt{CLS}]$ as a global token for classification. As in Tab. 6, we find that without $[\mathtt{CLS}]$ and utilizing only the average pooling of the final block’s output yields the best results for PointMamba.
@@
Analysis on dimension $C$. The dimension $C$ for the Mamba block is a crucial hyper-parameter for controlling the model’s scale. As in Tab. 7, we empirically set $C$ to 384 by considering the trade-off in performance across three variants.